OpenStudy (anonymous):

More cards! Consider again a standard deck of 52 cards (13 in each of 4 suits). Two cards are dealt in succession (meaning no replacement). a) What is the probability that both cards are greater than 7 (assuming that the ace is considered “high” or greater than 7)? b)Given that the first card was a queen, what is the probability that the second card is a seven? c) What is the probability of being dealt a queen first followed by a seven? d) What is the probability that both cards are greater than 7 (assuming that the ace is considered “high” or greater than 7)?

OpenStudy (anonymous):

how many cards are greater than 7?

OpenStudy (anonymous):

i count \(7\times 4=28\) so the answer to the first one is \[ \frac{28}{52}\times \frac{27}{51}\] whatever that is

OpenStudy (anonymous):

Given that the first card was a queen, what is the probability that the second card is a seven the first card is a queen there are 4 sevens left, and 51 cards left \[\frac{4}{51}\] for that one

OpenStudy (anonymous):

So for the first one, would be 756/2652 ?

OpenStudy (anonymous):

i don't know, i didn't do it

OpenStudy (anonymous):

cancel first, multiply last

OpenStudy (anonymous):

so 63/221 ?

OpenStudy (anonymous):

that is what the wolf says, yes

OpenStudy (anonymous):

okay so how would i do C now?

OpenStudy (anonymous):

first card is a queen what is the probability of that?

OpenStudy (anonymous):

4/52 ?

OpenStudy (anonymous):

is B and C asking the same thing?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i mean yes, \(\frac{4}{52}\) is right, but B and C are different questions

OpenStudy (anonymous):

This is was i have so far a) 63/221 b) 4/51 c) d)

OpenStudy (anonymous):

\(\frac{4}{52}\) is the probability that the first one is a queen \(\frac{4}{51}\) is the probability that the second one is a 7 IF the first one is a queen (that is why the denominator is 51 instead of 52)

OpenStudy (anonymous):

to find your answer, multiply \[\frac{4}{52}\times \frac{4}{51}\]

OpenStudy (anonymous):

A and D are identical questions

OpenStudy (anonymous):

1/663 for c???

OpenStudy (anonymous):

im confused on C

OpenStudy (anonymous):

lets go slow

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

the probability that the first card is a queen is \[\frac{4}{52}=\frac{1}{13}\]

OpenStudy (anonymous):

now you need the probability that the second card is a 7, given that the first card was a queen because one queen is out of the deck, there are 51 cards left of those 51 cards, 4 are sevens that probability is \[\frac{4}{51}\]

OpenStudy (anonymous):

yes so that is my answer to b? what is the difference in questions between b and c

OpenStudy (anonymous):

to get the probability that the first is a queen AND the second is a seven, you multiply those two probabilities, i.e. compute \[\frac{1}{13}\times \frac{4}{51}\]

OpenStudy (anonymous):

that was the answer to C, not to B

OpenStudy (anonymous):

so 4/663

OpenStudy (anonymous):

ok the purpose of question B, whose answer is \(\frac{4}{51}\) was to make you realize that the answer to C is \(\frac{1}{13}\times \frac{4}{51}\)

OpenStudy (anonymous):

okay phew thank you, so this is what i have now: a) 63/221 b) 4/51 c) 4/663 d) 63/221

OpenStudy (anonymous):

unless there is a typo somewheres, A and D are identical questions they look identical to you?

OpenStudy (anonymous):

yes they look identical to me too, so i just put the same answer

OpenStudy (anonymous):

thank you for helping me

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

The A.P. Statistics exam consists of a multiple-choice section where each question has five choices. Sara estimates that she has a 75% chance of knowing each question. If she does not know the answer, she will randomly guess. a) What is the probability of Sara getting a question correct? b) The A.P. exam has 40 questions, how many might Sara expect to get correct?

OpenStudy (anonymous):

any idea with this one?

OpenStudy (anonymous):

@satellite73